John plans to buy a vacation home in 5 years from now and wants to have saved $39,868 for a down payment. How much money should he place today in a saving account that earns 9.45 percent per year (compounded daily) to accumulate money for his down payment? Round the answer to two decimal places
Answer #1

Annual Interest Rate = 9.45 %, Daily Interest Rate = 9.45 / 365 ~ 0.02589 %

Target Future Value = $ 39868 and Tenure = 5 years or (5 x 365) = 1825 days

Let the initial deposit value be $ K

Therefore, K x (1.0002589)^(1825) = 39868

K = 39868 / (1.0002589)^(1825) = $ 24856.89